OpenStudy (galvan100):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1)

OpenStudy (mathmate):

p^2=1/2500, p=1/50 q=1-p=49/50 Heterozygous carriers are either pq or qp, so 2pq=2*49/50*1/50=98/2500